If the displacement is in the form of x=3/2(t)(2)+t+1 find the velocity at time 1 second

Grade:11

2 Answers

Shreya Gautam
5 years ago
Dear student, your question is not so clear to me but I take it as x=(3/2)t^2+t+1As we know, v is differentiation of x with respect to t, therefore we differentiate the given expression with respect to t. v= dx/dt= 3t +1. Therefore velocity at t=1 is 3*1+1= 4
Kaviya
5 years ago
If X=3/2(t)(2)+t+1
Now this is displacement, we should change that equation into velocity, which is
By differentiation, dx/dt = 3t+1
Now put the value of t=1 in the above equation
3(1)+1= 4m/s